# PoseAdapt: Sustainable Human Pose Estimation via Continual Learning Benchmarks and Toolkit

### v2

@@ 1 Introduction
Human pose estimation enables applications across autonomous systems [guesdon2021dripe, zheng2022multi, wang2023learning], healthcare [dibenedetto2024new, zhou2024gait, yin2025progait], sports analytics [ingwersen2023sportspose, yeung2025athletepose3d], and human–computer interaction [anvari20223d, huo20233d, son2025interaction]. Advances in representation methods [zhang2020distribution, wang2022contextual, li2022simcc, purkrabek2025probpose], model architectures [maji2022yolo, xu2022vitpose, jiang2023rtmpose], data scale [andriluka20142d, liu2020new, moon2020interhand, jin2020whole, khan2025towards], and training pipelines [joo2021exemplar, marusov2022enhancing] have pushed 2D keypoint benchmarks toward saturation [gao2025systematic].
@@
Despite these gains, state‑of‑the‑art models are fundamentally static: they are trained once on fixed datasets and deployed under the assumption that test distributions match training. In practice, changes in illumination, viewpoint, density, or sensing modality cause significant drops in accuracy [essich2023auxiliary, chellappa2024some]. These limitations are especially visible in dynamic or resource‑constrained settings (e.g., high‑speed sports [yeung2025athletepose3d, ingwersen2023sportspose] or egocentric capture [wang2022estimating, huang2024egoexolearn]), where strong motion, occlusion, and field‑of‑view shifts challenge standard training regimes. These domains often involve domain-specific skeletons or sensing modalities for which pretrained estimators are either mismatched or incomplete. As shown in Fig.˜2, off‑the‑shelf models degrade under realistic shifts, undermining their deployment-time performance despite high accuracy on benchmark datasets.
To overcome these deployment-time limitations or skeleton mismatches, off-the-shelf models often need to be fine-tuned on domain-specific data before they can be integrated in a real-world application. In Fig.˜3, we compare different adaptation strategies used in practice. Retraining a model from scratch for each deployment condition is costly and slow, and naive fine‑tuning tends to overwrite prior knowledge, leading to catastrophic forgetting [kirkpatrick2017overcoming, li2017learning]. Some recent efforts target cross‑skeleton generalization [chen2025recurrent], but typically rely on large backbones or extensive supervision, limiting deployability on edge devices. In contrast, continual learning (CL) methods regularize updates to preserve previously acquired knowledge while specializing to new experiences [kirkpatrick2017overcoming, li2017learning]. Therefore, we advocate continual adaptation as a sustainable alternative. Instead of discarding prior competence, models should incrementally incorporate new domains or keypoints while retaining past performance.
@@ 2 Related Work
Pose Estimation Toolkits and Distribution Shifts. Open-source systems such as OpenPose [cao2019openpose], Detectron2 [wu2019detectron2], MMPose [mmpose2020], AlphaPose [fang2022alphapose], RTMPose [jiang2023rtmpose], and ViTPose [xu2022vitpose] have driven rapid progress in 2D human pose estimation with modular training and evaluation across datasets and backbones. Despite their breadth, these toolkits largely assume a static train–test regime in which models are trained once on curated data and deployed unchanged. Related adaptation efforts in vision—for example, test-time adaptation via entropy minimization [wang2021tent]—offer robustness within a domain but do not address incremental task streams, skeleton growth, or long-horizon retention. Some cross-skeleton methods jointly train on multiple datasets [sarandi2023learning], which is inefficient and impractical for continual evolution. Consequently, deploying pose estimators in dynamic settings (wearables, robotics, sports analytics) remains brittle when illumination, viewpoint, density, or sensing modality shift over time. PoseAdapt framework provides continual adaptation protocols for pose estimation, missing in existing systems.
Benchmarks and Datasets for Pose Estimation. Large-scale benchmarks (COCO Keypoints [lin2014microsoft], MPII [andriluka20142d]) and newer efforts (COCO-WholeBody [jin2020whole], PoseTrack [andriluka2018posetrack], Halpe [fang2022alphapose], SpineTrack [khan2025towards]) have standardized evaluation and spurred architectural advances. However, these benchmarks are organized around static splits and assume full retraining when the task changes (e.g., skeleton extensions or cross-modality transfer). In contrast, continual learning (CL) benchmarks in classification study nonstationary streams and forgetting using protocols such as Split-CIFAR [zenke2017continual], CORe50 [lomonaco2017core50], and DomainNet [peng2019moment]. An analogous, pose-specific benchmark that enforces realistic constraints (fixed lightweight backbones, no access to past data, limited budgets) has been lacking; PoseAdapt fills this gap with domain- and class-incremental tracks.
@@
Continual Learning: Paradigms and Constraints. CL aims to learn from nonstationary streams without catastrophic forgetting [mccloskey1989catastrophic, parisi2019continual, van2022three]. Representative approaches include: regularization-based methods that constrain parameter or function drift (EWC [kirkpatrick2017overcoming], SI [zenke2017continual], LwF [li2017learning], LFL [jung2016less], IMM [lee2017overcoming]); replay-based methods that use exemplars or generative rehearsal (iCaRL [rebuffi2017icarl], DGR [shin2017continual], A-GEM [chaudhry2019efficient], DER [buzzega2020dark]); and architectural approaches that isolate or grow capacity (PNNs [rusu2016progressive], PackNet [mallya2018packnet], Piggyback [mallya2018piggyback], HAT [serra2018overcoming]). Many of these assume either access to a replay buffer or the ability to expand parameters and heads per task, which conflicts with deployment constraints such as fixed memory, strict compute, and privacy limits on past data. PoseAdapt benchmark isolates adaptation strategy effects by fixing model architecture, disallowing access to old data, and enforcing per-step budgets, enabling fair comparison under realistic constraints.
Continual Learning for Dense Prediction. Transferring CL to dense prediction introduces unique challenges: spatial outputs, structured losses, and label-evolution (e.g., new classes or keypoints). Recent work in segmentation [cermelli2020modeling, michieli2019incremental] and detection [shmelkov2017incremental] adapts regularization, distillation, or pseudo-labeling to retain prior categories while learning new ones. For human pose estimation, CL remains comparatively underexplored; early studies adapt CL ideas to structured outputs [gu2022not], but evaluations are heterogeneous, often using small-scale or ad hoc protocols. By standardizing domain-incremental and class-incremental settings with shared budgets and metrics (forward transfer, retention), PoseAdapt provides the first controlled testbed to assess CL strategies for pose at scale.
Beyond benchmarks, software ecosystems for CL such as Avalanche [lomonaco2021avalanche] and Continuum [douillard2021continuum] provide abstractions and baselines, while general domain-shift suites (e.g., WILDS [koh2021wilds]) and domain generalization toolkits (e.g., DomainBed [gulrajani2021in]) target classification or detection tasks. MMPose and Detectron2 offer strong pose tooling but assume static training. PoseAdapt complements these ecosystems by bringing pose-specific continual protocols, plug-in CL strategies, and unified evaluation under deployment-motivated constraints, enabling reproducible comparisons and practical adaptation pathways.
@@ 4.1 Experimental Setup
Methods. We evaluate naïve fine-tuning (FT), Elastic Weight Consolidation (EWC) [kirkpatrick2017overcoming], Less-Forgetful Learning (LFL) [jung2016less], and Learning without Forgetting (LwF) [li2017learning].
@@
Reference model and data. All methods adapt the same off-the-shelf top-down pose estimator. To isolate keypoint regression from person detection, ground‑truth detection boxes are used. The model (RTMPose‑t [jiang2023rtmpose], $\sim$3M parameters) is pretrained on the COCO [lin2014microsoft] and AIC datasets to predict 17 keypoints with 70.06 AP on the COCO validation set [^2] https://download.openmmlab.com/mmpose/v1/projects/rtmposev1/rtmpose-tiny_simcc-aic-coco_pt-aic-coco_420e-256x192-cfc8f33d_20230126.pth. We denote this as the reference domain (Fig.˜5) in the first experience which the model must not forget as it adapts to new experiences with various domain shifts.
Metrics. The mean average precision (AP) at the end of training all experiences is reported. In addition, following standard continual learning evaluation [lopez2017gradient], we report retention accuracy (RA) and average forgetting (AF). Let $a_{i,j}$ denote the AP on experience $j$ after training on experience $i$. For a sequence of $T$ experiences:
@@
Hyperparameters and constraints. Adaptation uses AdamW [loshchilov2017decoupled] with a base learning rate 0.004, no weight decay, random seed 21, and a small linear warm-up. All images are normalized with $\mu=[123.675,116.28,103.53],\sigma=[58.395,57.12,57.375]$ and augmented using random horizontal flips and half-body occlusions. Each experience is allocated at most 1k labeled images and 10 epochs. Past data are never stored or revisited. The only state retained across experiences is the most recent network; distillation‑based methods additionally keep a teacher snapshot from the previous experience. For EWC, the diagonal Fisher and parameter means are computed at the end of each experience on its training stream and stored for the next step.
@@ 4.2.3 Modality
Shift generation. Modality shifts comprise two experiences: Gray and Depth. Grayscale images are produced by fully desaturating each RGB frame and then applying a sequence of mild photometric perturbations including some Gaussian noise, increased contrast, and a small brightness correction. Depth images are generated by running MiDaS (DPT-Large) [ranftl2020midas] per frame, resizing predictions back to the original resolution, and min–max normalizing the relative-depth map to [0,1]. Both outputs, illustrated in Fig.˜8, are tiled to three channels for backbone compatibility.
@@ 4.3 Class‑Incremental Track
For standard evaluations of pure class-incremental continual pose estimation, we propose the PoseAdapt-BodyParts benchmark. In this benchmark, the network must progressively learn body, feet, hands, face, and spine keypoints. Body annotations with 17 keypoints from the COCO dataset [lin2014microsoft], feet, face, and hand annotations with 6, 68, and 21 per hand keypoints respectively from the COCO-Wholebody dataset [jin2020whole], and 9 spine annotations from SpineTrack [khan2025towards] are used. Across the five experiences, the number of keypoints changes from 17, 23, 91, and 133 to 142 in the last experience. Training images from COCO dataset, shared by all three datasets, rule out any domain shifts. This allows the benchmark to solely assess skeleton growth capabilities of different continual learning strategies. As in the domain-incremental track, an off-the-shelf body pose estimator is used for the first experience, and the other four experiences are fine-tuned with maximum 10 epochs per experience.
@@ 5 Conclusion
Acknowledgments. This work has received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101092889.

### v3

@@ 1 Introduction
Human pose estimation enables applications across autonomous systems [16, 67, 60], healthcare [10, 68, 64], sports analytics [20, 63], and human–computer interaction [3, 19, 55]. Advances in representation methods [66, 58, 29, 47], model architectures [37, 62, 21], data scale [2, 32, 44, 22, 25], and training pipelines [23, 40] have pushed 2D keypoint benchmarks toward saturation [14].
@@
Despite these gains, state‑of‑the‑art models are fundamentally static: they are trained once on fixed datasets and deployed under the assumption that test distributions match training. In practice, changes in illumination, viewpoint, density, or sensing modality cause significant drops in accuracy [12, 8]. These limitations are especially visible in dynamic or resource‑constrained settings (e.g., high‑speed sports [63, 20] or egocentric capture [59, 18]), where strong motion, occlusion, and field‑of‑view shifts challenge standard training regimes. These domains often involve domain-specific skeletons or sensing modalities for which pretrained estimators are either mismatched or incomplete. As shown in Fig.˜2, off‑the‑shelf models degrade under realistic shifts, undermining their deployment-time performance despite high accuracy on benchmark datasets.
To overcome these deployment-time limitations or skeleton mismatches, off-the-shelf models often need to be fine-tuned on domain-specific data before they can be integrated in a real-world application. In Fig.˜3, we compare different adaptation strategies used in practice. Retraining a model from scratch for each deployment condition is costly and slow, and naive fine‑tuning tends to overwrite prior knowledge, leading to catastrophic forgetting [26, 30]. Some recent efforts target cross‑skeleton generalization [9], but typically rely on large backbones or extensive supervision, limiting deployability on edge devices. In contrast, continual learning (CL) methods regularize updates to preserve previously acquired knowledge while specializing to new experiences [26, 30]. Therefore, we advocate continual adaptation as a sustainable alternative. Instead of discarding prior competence, models should incrementally incorporate new domains or keypoints while retaining past performance.
@@ 2 Related Work
Pose Estimation Toolkits and Distribution Shifts. Open-source systems such as OpenPose [5], Detectron2 [61], MMPose [43], AlphaPose [13], RTMPose [21], and ViTPose [62] have driven rapid progress in 2D human pose estimation with modular training and evaluation across datasets and backbones. Despite their breadth, these toolkits largely assume a static train–test regime in which models are trained once on curated data and deployed unchanged. Related adaptation efforts in vision—for example, test-time adaptation via entropy minimization [57]—offer robustness within a domain but do not address incremental task streams, skeleton growth, or long-horizon retention. Some cross-skeleton methods jointly train on multiple datasets [51], which is inefficient and impractical for continual evolution. Consequently, deploying pose estimators in dynamic settings (wearables, robotics, sports analytics) remains brittle when illumination, viewpoint, density, or sensing modality shift over time. PoseAdapt framework provides continual adaptation protocols for pose estimation, missing in existing systems.
Benchmarks and Datasets for Pose Estimation. Large-scale benchmarks (COCO Keypoints [31], MPII [2]) and newer efforts (COCO-WholeBody [22], PoseTrack [1], Halpe [13], SpineTrack [25]) have standardized evaluation and spurred architectural advances. However, these benchmarks are organized around static splits and assume full retraining when the task changes (e.g., skeleton extensions or cross-modality transfer). In contrast, continual learning (CL) benchmarks in classification study nonstationary streams and forgetting using protocols such as Split-CIFAR [65], CORe50 [33], and DomainNet [46]. An analogous, pose-specific benchmark that enforces realistic constraints (fixed lightweight backbones, no access to past data, limited budgets) has been lacking; PoseAdapt fills this gap with domain- and class-incremental tracks.
@@
Continual Learning: Paradigms and Constraints. CL aims to learn from nonstationary streams without catastrophic forgetting [41, 45, 56]. Representative approaches include: regularization-based methods that constrain parameter or function drift (EWC [26], SI [65], LwF [30], LFL [24], IMM [28]); replay-based methods that use exemplars or generative rehearsal (iCaRL [49], DGR [53], A-GEM [7], DER [4]); and architectural approaches that isolate or grow capacity (PNNs [50], PackNet [39], Piggyback [38], HAT [52]). Many of these assume either access to a replay buffer or the ability to expand parameters and heads per task, which conflicts with deployment constraints such as fixed memory, strict compute, and privacy limits on past data. PoseAdapt benchmark isolates adaptation strategy effects by fixing model architecture, disallowing access to old data, and enforcing per-step budgets, enabling fair comparison under realistic constraints.
Continual Learning for Dense Prediction. Transferring CL to dense prediction introduces unique challenges: spatial outputs, structured losses, and label-evolution (e.g., new classes or keypoints). Recent work in segmentation [6, 42] and detection [54] adapts regularization, distillation, or pseudo-labeling to retain prior categories while learning new ones. For human pose estimation, CL remains comparatively underexplored; early studies adapt CL ideas to structured outputs [15], but evaluations are heterogeneous, often using small-scale or ad hoc protocols. By standardizing domain-incremental and class-incremental settings with shared budgets and metrics (forward transfer, retention), PoseAdapt provides the first controlled testbed to assess CL strategies for pose at scale.
Beyond benchmarks, software ecosystems for CL such as Avalanche [34] and Continuum [11] provide abstractions and baselines, while general domain-shift suites (e.g., WILDS [27]) and domain generalization toolkits (e.g., DomainBed [17]) target classification or detection tasks. MMPose and Detectron2 offer strong pose tooling but assume static training. PoseAdapt complements these ecosystems by bringing pose-specific continual protocols, plug-in CL strategies, and unified evaluation under deployment-motivated constraints, enabling reproducible comparisons and practical adaptation pathways.
@@ 4.1 Experimental Setup
Methods. We evaluate naïve fine-tuning (FT), Elastic Weight Consolidation (EWC) [26], Less-Forgetful Learning (LFL) [24], and Learning without Forgetting (LwF) [30].
@@
Reference model and data. All methods adapt the same off-the-shelf top-down pose estimator. To isolate keypoint regression from person detection, ground‑truth detection boxes are used. The model (RTMPose‑t [21], $\sim$3M parameters) is pretrained on the COCO [31] and AIC datasets to predict 17 keypoints with 70.06 AP on the COCO validation set [^2] https://download.openmmlab.com/mmpose/v1/projects/rtmposev1/rtmpose-tiny_simcc-aic-coco_pt-aic-coco_420e-256x192-cfc8f33d_20230126.pth. We denote this as the reference domain (Fig.˜5) in the first experience which the model must not forget as it adapts to new experiences with various domain shifts.
Metrics. The mean average precision (AP) at the end of training all experiences is reported. In addition, following standard continual learning evaluation [35], we report retention accuracy (RA) and average forgetting (AF). Let $a_{i,j}$ denote the AP on experience $j$ after training on experience $i$. For a sequence of $T$ experiences:
@@
Hyperparameters and constraints. Adaptation uses AdamW [36] with a base learning rate 0.004, no weight decay, random seed 21, and a small linear warm-up. All images are normalized with $\mu=[123.675,116.28,103.53],\sigma=[58.395,57.12,57.375]$ and augmented using random horizontal flips and half-body occlusions. Each experience is allocated at most 1k labeled images and 10 epochs. Past data are never stored or revisited. The only state retained across experiences is the most recent network; distillation‑based methods additionally keep a teacher snapshot from the previous experience. For EWC, the diagonal Fisher and parameter means are computed at the end of each experience on its training stream and stored for the next step.
@@ 4.2.3 Modality
Shift generation. Modality shifts comprise two experiences: Gray and Depth. Grayscale images are produced by fully desaturating each RGB frame and then applying a sequence of mild photometric perturbations including some Gaussian noise, increased contrast, and a small brightness correction. Depth images are generated by running MiDaS (DPT-Large) [48] per frame, resizing predictions back to the original resolution, and min–max normalizing the relative-depth map to [0,1]. Both outputs, illustrated in Fig.˜8, are tiled to three channels for backbone compatibility.
@@ 4.3 Class‑Incremental Track
For standard evaluations of pure class-incremental continual pose estimation, we propose the PoseAdapt-BodyParts benchmark. In this benchmark, the network must progressively learn body, feet, hands, face, and spine keypoints. Body annotations with 17 keypoints from the COCO dataset [31], feet, face, and hand annotations with 6, 68, and 21 per hand keypoints respectively from the COCO-Wholebody dataset [22], and 9 spine annotations from SpineTrack [25] are used. Across the five experiences, the number of keypoints changes from 17, 23, 91, and 133 to 142 in the last experience. Training images from COCO dataset, shared by all three datasets, rule out any domain shifts. This allows the benchmark to solely assess skeleton growth capabilities of different continual learning strategies. As in the domain-incremental track, an off-the-shelf body pose estimator is used for the first experience, and the other four experiences are fine-tuned with maximum 10 epochs per experience.
@@ 5 Conclusion
Acknowledgement.
The work leading to this publication was co-funded by the European Union’s Horizon Europe research and innovation programme under Grant Agreement No 101135724 (project LUMINOUS) and Grant Agreement No 101092889 (project SHARESPACE).